### v1

@@ I Introduction
Quantum computers may outperform classical computers in a variety of tasks [1, 2, 3], but these are so far inaccessible; despite progress in realising these devices across a variety of physical platforms [4, 5, 6, 7], building and running quantum computers is associated with a large financial cost [8, 9, 10]. Classical cloud-based access to quantum servers eliminates the need for users to own large and expensive devices themselves [11], but is unsuitable for use cases involving sensitive data, in which a user requires access to the computational power of the quantum server without revealing the input data, the computation or the output to the owner of the server [12, 13, 14].
@@
Blind quantum computing (BQC) is a technique where a client can perform quantum algorithms at a remote quantum server without the input, the computation, or its outcome being revealed (apart from an upper bound on the size of the computation) [15]. Preferably, the client is realised as cheaply as possible, to help make quantum computing more widely available. Two options for realising such client are for it to have the ability to either send single photons, or measure them [16]. In this work, we assume quantum capabilities of the client to be limited to making measurements.
@@
The initial BQC protocol, based on measurement-based quantum computation [17, 18], was later expanded to include verification to test for correctness [19]. In such verifiable blind quantum computing (VBQC) protocols, the client can abort if the output of certain tests are not as expected.
@@
In realistic near-term noisy quantum devices, imperfections are inevitable. In verifying tests, imperfections and noise in the system can be mistaken for malicious behaviour of the server, resulting in a computation that will be aborted constantly, and the client gaining little information about the operations of the server. To combat this, a noise-robust verified BQC (rVBQC) protocol was introduced [20]. This protocol tolerates imperfections from noise or malicious behaviour provided that the server does not fail more than 25% of verifying tests employed by the client (for the type of computation we study, see Section III for details). The robustness to noise is realised by repeating the computation multiple times and performing classical error correction (majority voting) on the results.
Already, there have been proof-of-principle demonstrations of BQC [21] and rVBQC [22] in laboratory settings. For real-world practicality, however, the client needs to be able to be spatially remote from the server, which will require improvements to existing quantum computing and communication hardware.
@@
In this paper, we determine the requirements for performing rVBQC at a metropolitan scale (i.e., the scale of a large city) of 50 km using a trapped-ion based server and a measurement-only client as depicted in Figure 1, using the 25% error tolerance as a threshold. In measurement-based quantum computing, any single-qubit gate can be performed using a five-qubit linear graph state (i.e., five qubits, each in a superposition state, with controlled-$Z$ operations between them) [23]. We use this five-qubit graph state to benchmark the performance of the protocol. We focus on rVBQC because this is feasibly achievable near-term.
@@
We investigate this implementation of the rVQBC protocol numerically using NetSquid, a discrete event simulator for quantum networks [24]. To this end, we introduce a framework for the modelling of trapped-ion quantum servers, including a NetSquid library [25], along with a model of a measurement-only client. The simulation is hardware motivated and takes a set of hardware parameters as input. Using this, we both:
@@
    Identify the per-parameter minimal requirements for hardware to allow for rVBQC. In each case, we assume perfect performance of all other parameters apart from fibre attenuation. This gives us a strict lower-bound for each parameter, which is compared to state-of-the-art performance. We show that current state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis, but all current imperfections combined make it impossible to perform rVBQC securely over 50 km using existing technology. These results are summarised in Figure 4;
@@
    Identify the set of hardware parameters that minimises the total improvement needed over current state-of-the-art parameters to allow for a successful implementation of rVBQC. This reveals which parameters need the most improvement and how far we need to improve them to enable a metropolitan-scale application of rVBQC. We do this by combining our requirement on the error probability as well as the cost ([26, 27]) of a set of hardware parameters into a single-objective minimisation problem. This is fed into a genetic algorithm [28] that minimises the total improvement needed over state-of-the-art performance. These results show there is substantial work left to be done in hardware improvements, and are summarised in Figure 5.
We organise this paper as follows. In Section II, we provide details of the physical setup that we simulate, including physical parameters. Then in Section III, we analyse the two primary results of our numerical analysis discussed briefly above. Section IV details the trapped-ion model, the client’s measurement apparatus, as well as the genetic algorithm and cost function that define our optimisation procedure. Finally, in Section V we discuss possible directions for future work.
@@ II Setup
We simulate a two-party rVBQC setup with a trapped-ion quantum server and a measurement-only client using NetSquid. We investigate the protocol at a metropolitan scale, in which the server and client are separated by 50 km of optical fibre. An overview of the setup is provided in Figure 1. The protocol used here, as in [22], assumes a variation wherein the client uses measurements to perform remote state preparation (RSP) on the server. In RSP, a sender measures part of an entangled state and communicates a classical correction to prepare a target state at a receiver. A formal description of the general protocol can be found in A.
Below, we will outline how normally rVBQC includes both computation and test rounds, and why our analysis only focuses on test rounds (Section II.1). In Sections II.2 through II.5 we describe the steps of such a test round. A visualisation of these steps is given in Figure 2. After this, we introduce the parameter sets that are used as input to the simulation (Section II.6) and the metrics on which we base our analysis (Section II.7). This provides necessary context to understand the results as presented in Section III.
@@ II.1 Computation and test rounds
In a normal run of the protocol, one picks a total number of rounds $N$ that are separated into computation rounds and test rounds, as suggested in [20]. The output of the total computation is taken as the majority output of the computation rounds.
For a test round we make use of trap qubits that are prepared in a superposition basis and are decoupled from the rest of the graph by having them surrounded by dummy qubits. Dummy qubits are prepared in the standard basis. The client asks the server to measure the traps in the basis they are prepared in and is therefore aware of what the outcome should be. A test round fails when any of the measurement outcomes of the traps do not equal the expected outcome. A test can fail due to noise or due to malicious behaviour by the server. In the computation considered here, the error tolerance for test rounds is 25%, as is explained in Section IV.3.
In our analysis, we focus only on test rounds as the computation rounds do not provide additional information about the verifiability of the setup. From the test rounds, we extract the error probability as the fraction of failed rounds. With the resource we consider, a five-qubit linear graph (i.e., five qubits connected through $CZ$ gates)— the test rounds can be accompanied with computation rounds in which any single-qubit gate is performed [29]. Thus, our results are universal for single-qubit gates.
We will outline how the test rounds are performed by going through the steps in Sections II.2 through II.5, Figure 2 provides an overview of these steps.
@@ II.2 Remote state preparation
The client controls the state of the qubits at the server by performing remote state preparation (RSP). In this work, we assume the quantum capabilities of the client to be limited to making measurements. Using entangled qubit pairs, the client can project a qubit located in the server onto a chosen basis by measuring an entangled qubit emitted (and sent to the client). In this way, the state of the qubits is known to (and determined by) the client, but it is unknown to the server. This is the source of the blindness of this protocol.
@@
In each round, the client randomly chooses the sets of qubits that will become the trap qubits and that will become the dummy qubits, such that all traps are surrounded by dummies. For a five-node graph, one can have either dummy-trap-dummy-trap-dummy or trap-dummy-trap-dummy-trap (other combinations are sub-optimal for trap insertion). The dummy qubits need to be prepared in the standard basis $H/V$ (i.e., on the north or south pole of the Bloch sphere) and the trap qubits need to be prepared in a superposition basis |±$\ket{\pm_{\theta_{v}}}=(\ket{H}\pm e^{i\theta_{v}}\ket{V})/\sqrt{2}$ (i.e., on the equator of the Bloch sphere) where $\theta_{v}\in\{k\pi/4\}_{0\leq k\leq 7}$, with $k$ an integer such that the angle $\theta_{v}$ is randomly and independently chosen for each qubit $v$.
@@
To start off RSP, the client communicates a description of the graph it wants to prepare to the server, this graph is the same for all rounds. This description includes the nodes in the graph $I$ and the edges of the graph $G$ (describing which nodes are connected through $CZ$ gates). Here, we consider a five-qubit linear graph ($I=\{0,1,2,3,4\};G=\{(0,1),(1,2),(2,3),(3,4)\}$).
@@
After receiving the description, the server emits a polarisation-encoded photonic qubit that is entangled with an ion as $|\psi\rangle=(\ket{0H}+\ket{1V})/\sqrt{2}$ and this photonic qubit is sent to the client over 50 km of optical fibre. The client then measures this photon in the standard or superposition basis, depending on whether it is a trap or dummy. The client thereby remotely prepares five qubits on the server that are in the state |$\ket{m_{v}}$ or |$\ket{+_{\theta_{v}^{\prime}}}=(\ket{0}+e^{i\theta_{v}^{\prime}}\ket{1})/\sqrt{2}$, where $\theta_{v}^{\prime}=\theta_{v}+m_{v}\pi$, with $m_{v}$ the outcome of the client’s measurement. Note that the server does not know the basis in which the client measures, and therefore it is unaware of the state of the qubits in its memory. Because the server does not know the state of the qubits, it also cannot tell the difference between a computation round (which involves only |±$\ket{\pm_{\theta_{v}}}$-qubits) and a test round (which includes dummy qubits) without performing a (malicious) intermediate measurement that has a chance of disturbing the quantum state.
@@
The client performs the superposition basis measurement by rotating the polarisation state of the incoming ion-entangled-photonic qubit followed by measuring the qubit in the standard basis. Rotating the polarisation state of the photonic qubit can be done by optical elements such as waveplates or electro-optic modulators. A measurement in the standard basis can be performed by a polarising beam splitter (PBS) followed by single photon detectors (SPD). For a more detailed description of a possible physical realisation of the client, see [22]. The simulated rotation setup consists of one half waveplate, followed by two quarter waveplates based on [30].
@@ II.3 Cutoff time
In the RSP process, some qubits will have been prepared and will be sitting in the memory, waiting for the preparation of the remaining qubits. Depending on the efficiency of the server and due to loss from the fibre, remotely preparing a qubit might take many tries, meaning that the qubits may have to reside in the memory for a long time. The longer the qubits are in memory, the more they decohere (details in Section IV.1). For a lower error probability, we want to limit the time qubits spend in memory, for which we include a cutoff time. This cutoff ensures that no qubit resides in the server memory for longer than half a coherence time during the RSP phase, thereby limiting the decoherence.
Choosing a good cutoff time is a trade-off between rate and fidelity: a shorter cutoff time gives higher quality qubits as they have suffered from less decoherence, yet more qubits will be discarded, leading to a lower rate. However, since this work only targets an error probability (which depends on the fidelity of the remotely prepared qubits) and does not consider rate, no optimisation can be performed on this aspect. See e.g. [31] for examples of optimising this. Instead, we choose a fixed cutoff of half the coherence time.
When any of the qubits in memory have been at the server for longer than the cutoff time, they are discarded and the client is notified. When the server has prepared five qubits, no more qubits will be discarded (i.e., no more cutoff after RSP).
@@ II.4 Graph formation
The client instructs the server to arrange the qubits in the linear graph: edges should be made between qubits $(0,1),(1,2),(2,3)$ and $(3,4)$. The server does this by applying controlled-$Z$ ($CZ$) operations between successive qubits to form a chain-like graph.
We assume that the ion qubits can be individually addressed for rotations $R_{x},R_{y},R_{z}$. This ability is hardware-dependent (it is possible in, for example, [32, 33, 34]). Other options besides direct-addressing include global addressing of ions either by ‘hiding’ other ions (as in [35]), ‘shuttling’ ions to different physical regions (as in [36, 37, 38]) or adding extra rotations (as in [39]), as long as rotations around at least one axis can be individually addressed.
We assume the bichromatic gate $R_{XX}(\theta)=\exp(-i\theta/2\sum_{i<j}\sigma_{x}^{(i)}\sigma_{x}^{(j)})$, also known as the Mølmer-Sørensen gate [40, 41] to be the native entangling operation. We use Qiskit [42] to find how to construct a $CZ$ gate using the operations available for the ion trap and we show this in Figure 3.
In principle, the server does not need to wait for all qubits to be prepared before it starts applying the CZ operations, but it is done so in the simulation for simplicity. The time it takes to prepare the graph state (around 150 $\mu$s) is short compared to the expected time it takes for the qubits to be remotely prepared (in the order of tens of milliseconds  [^1] Calculated as time it takes to perform one attempt (photon round-trip time: 2x50 km / (c/n) $\approx$ 0.50 ms, with c the speed of light and n = 1.4682 the index of refection of telecommunication optical fibre) divided by the average probability of photon arriving at client (server efficiency times probability of photon transmitting through the fibre: 0.1325 $\times$ (50km $\times$ 0.2 dB/km)=0.1. Giving an average probability of photon arriving of 0.01325) giving 0.50/0.01325=37.8 ms.).
@@ II.5 Server measurements and output
Once the graph state formation is complete, the client instructs the server to measure in ±$\pm_{\delta_{v}}$-bases defined by
@@
Here, $\theta_{v}$ defines the basis used by the client to perform RSP and $m_{v}$ is the outcome of the RSP measurement of the photonic qubit entangled with ion qubit $v$. $r_{v}$ is a random bit generated by the client to ensure the measurement outcomes appear completely random. The angles at which the dummy qubits are measured is irrelevant to the client; they just needs to be random, such that the server will not be able to detect the qubit as a dummy (which would reveal that the client is executing a test round).
The server communicates the outcome of measuring qubit $v$ to the client as $b_{v}$. For all trap qubits, the client checks if $b_{v}=r_{v}$. The test round succeeds if this is true for all traps. If any of the traps do not satisfy $b_{v}=r_{v}$, the round fails. The round failing or succeeding is the output for each of the test rounds.

### v2

@@ I Introduction
Quantum computers may outperform classical computers in a variety of tasks [1, 2, 3], but these advantages are so far inaccessible. Moreover, despite progress in realising these devices across a variety of physical platforms [4, 5, 6, 7, 8], building and running quantum computers is associated with a large financial cost [9, 10, 11]. Cloud-based access to quantum servers eliminates the need for users to own large and expensive devices themselves [12] but is unsuitable for use cases involving sensitive data, in which a user requires access to the computational power of the quantum server without revealing the input data, the computation or the output to the owner of the server [13, 14, 15].
@@
Blind quantum computing (BQC) is a technique with which a client can execute quantum algorithms at a remote quantum server without the input, the computation, or its outcome being revealed (apart from an upper bound on the size of the computation) [16]. Preferably, the client is realised as cheaply as possible, to help make quantum computing more widely available. Two options for realising such a client are for it to have the ability to either send single photons [16] or measure them [17, 18]. Alternatively, one can make use of a single-qubit-gate-performing client [19] or a multi-server approach, in combination with a completely classical client [20, 21, 22, 23]. In this work, we assume a single-server setup where the quantum capabilities of the client are limited to making measurements.
@@
The initial BQC protocol, based on measurement-based quantum computation [24, 25], was later expanded to include verification to test for correctness; approaches for verification are summarised in reference [26]. In such verifiable blind quantum computing (VBQC) protocols, the client can abort if the outputs of certain tests (either trap based [27], stabiliser based [28] or classical but introducing computational assumptions [29]) are not as expected.
@@
In realistic near-term noisy quantum devices [30], imperfections are inevitable. In verifying tests, imperfections and noise in the system can be mistaken for malicious behaviour of the server, resulting in a computation that will be aborted constantly and the client gaining little information about the operations of the server. There are ways to deal with noise in non-verified protocols [31], and later a noise-robust verified BQC (rVBQC) protocol was introduced aswell [32]. This rVBQC protocol tolerates imperfections from noise or malicious behaviour provided that the server does not fail more than 25% of verifying tests employed by the client. The robustness to noise is realised by repeating the computation multiple times and performing classical error correction (majority voting) on the results. Already, there have been proof-of-principle demonstrations of BQC [33, 34] and rVBQC [35] in laboratory settings. For real-world practicality, however, the client needs to be able to be spatially remote from the server, which will require improvements to existing quantum computing and communication hardware.
@@
In this paper, we determine the requirements for performing rVBQC at a metropolitan scale (i.e., the scale of a large city) of 50 km using a trapped-ion-based server and a measurement-only client as depicted in figure 1, using the 25% error tolerance as a threshold. In measurement-based quantum computing, any single-qubit gate can be performed using a five-qubit linear graph state (i.e., five qubits, each in a superposition state, with controlled-$Z$ operations between them) [36]. We use this five-qubit graph state to benchmark the performance of the protocol. We focus on rVBQC because this is feasibly achievable in the near-term.
@@
We investigate this implementation of the rVQBC protocol numerically using NetSquid, a discrete event simulator for quantum networks [37]. To this end, we introduce a framework for the modelling of trapped-ion quantum servers, including a NetSquid library [38], along with a model of a measurement-only client. The simulation is hardware motivated and takes a set of hardware parameters as input. Using this, we
@@
    Identify the per-parameter minimal requirements for hardware to allow for rVBQC. In each case, we assume perfect performance of all other parameters apart from fibre attenuation. This gives us a strict lower bound for each parameter, which is compared to state-of-the-art performance. We show that current state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis, but all current imperfections combined make it impossible to perform rVBQC securely over 50 km using existing technology. These results are summarised in figure 4;
@@
    Identify the set of hardware parameters that minimises the total improvement needed over current state-of-the-art parameters to allow for a successful implementation of rVBQC. This reveals which parameters need the most improvement and how far we need to improve them to enable a metropolitan-scale application of rVBQC. We do this by combining our requirement on the error probability with the cost ([39, 40]) of a set of hardware parameters into a single-objective minimisation problem. This is fed into a genetic algorithm [41] that minimises the total improvement needed over state-of-the-art performance. These results show there is substantial work left to be done in hardware improvements, and they are summarised in figure 5.
We organise this paper as follows. In section II, we provide details of the physical setup that we simulate, including physical parameters. Then in section III, we analyse the two primary results of our numerical analysis discussed briefly above. Section IV details the trapped-ion model, the client’s measurement apparatus, and the genetic algorithm and cost function that define our optimisation procedure. Finally, in section V we discuss possible directions for future work.
@@ II Setup
We simulate a two-party rVBQC setup with a trapped-ion quantum server and a measurement-only client using NetSquid. We investigate the protocol at a metropolitan scale, in which the server and client are separated by 50 km of optical fibre. An overview of the setup is provided in figure 1. The protocol used here, as in reference [35], assumes a variation wherein the client uses measurements to perform remote state preparation (RSP) on the server. In RSP, a sender measures part of an entangled state and communicates a classical correction to prepare a target state at a receiver.
Below, we will outline how normally rVBQC includes both computation and test rounds, and why our analysis only focuses on test rounds (section II.1). In sections II.2 through II.5 we describe the steps of such a test round, which summarises the protocol of reference [32] and how this is adapted for our simulations. A visualisation of these steps is given in figure 2. After this, we introduce the parameter sets that are used as input to the simulation (section II.6) and the metrics on which we base our analysis (section II.7). This provides necessary context to understand the results as presented in section III.
@@ II.1 Computation and test rounds
In a normal run of the protocol, one picks a total number of rounds $N$ that are separated into computation rounds and test rounds, as suggested in reference [32]. The output of the total computation is taken as the majority output of the computation rounds. A formal description of the general protocol can be found in A.
In our analysis, we focus only on test rounds as the computation rounds do not provide additional information about the verifiability of the setup. From the test rounds, we extract the error probability as the fraction of failed rounds. With the resource we consider, a five-qubit linear graph — the test rounds can be accompanied by computation rounds in which any single-qubit gate is performed [42]. Thus, our results are universal for single-qubit gates.
@@ II.2 Remote state preparation
The client controls the state of the qubits at the server via RSP. The client can project a qubit located at the server onto a chosen basis by measuring a second qubit, emitted by and entangled with the first qubit (and sent to the client). In this way, the state of the qubits is known to (and determined by) the client, but it is unknown to the server. This is the source of the blindness of this protocol.
@@
In each round, the client randomly chooses the sets of qubits that will become the trap qubits and the dummy qubits, such that all trap qubits are surrounded by dummies. For a five-node graph, one can have either dummy-trap-dummy-trap-dummy or trap-dummy-trap-dummy-trap (other combinations are sub-optimal for trap insertion). The dummy qubits need to be prepared in the standard basis $0/1$ (i.e., on the north or south pole of the Bloch sphere), and the trap qubits need to be prepared in a superposition basis |±$\ket{\pm_{\theta_{v}}}=(\ket{0}\pm e^{i\theta_{v}}\ket{1})/\sqrt{2}$ (i.e., on the equator of the Bloch sphere) where $\theta_{v}\in\{k\pi/4\}_{0\leq k\leq 7}$, with $k$ an integer such that the angle $\theta_{v}$ is randomly and independently chosen for each qubit $v$.
@@
To start off RSP, the client communicates a description of the graph it wants to prepare to the server. The graph description is the same for all rounds and does not reveal which of the qubits are trap qubits or that a test round is performed at all. This description includes the nodes in the graph $I$ and the edges of the graph $G$ (describing which nodes are connected through $CZ$ gates). Here, we consider a five-qubit linear graph ($I=\{0,1,2,3,4\};G=\{(0,1),(1,2),(2,3),(3,4)\}$).
@@
After receiving the description, the server emits a polarisation-encoded photonic qubit that is entangled with an ion as $|\psi\rangle=(\ket{0H}+\ket{1V})/\sqrt{2}$, and this photonic qubit is sent to the client over 50 km of optical fibre. The client then sends back a confirmation of arrival to the server and measures the photon in the standard or superposition basis, depending on whether it is a trap qubit or dummy. The client thereby remotely prepares five qubits on the server that are in the state |$\ket{m_{v}}$ or |$\ket{+_{\theta_{v}^{\prime}}}=(\ket{0}+e^{i\theta_{v}^{\prime}}\ket{1})/\sqrt{2}$, where $\theta_{v}^{\prime}=\theta_{v}+m_{v}\pi$, with $m_{v}$ the outcome of the client’s measurement. Note that the server does not know the basis in which the client measures, and therefore it is unaware of the state of the qubits in its memory. Because the server does not know the state of the qubits, it also cannot tell the difference between a computation round (which involves only |±$\ket{\pm_{\theta_{v}}}$-qubits) and a test round (which includes dummy qubits) without performing a (malicious) intermediate measurement that has a chance of disturbing the quantum state.
@@
The client performs the superposition basis measurement by rotating the polarisation state of the incoming ion-entangled photonic qubit, then measuring the qubit in the standard basis. Rotating the polarisation state of the photonic qubit can be done by optical elements such as waveplates or electro-optic modulators. A measurement in the standard basis can be performed by a polarising beam splitter (PBS) followed by single photon detectors (SPDs). For a more detailed description of a possible physical realisation of the client, see reference [35]. The simulated rotation setup consists of one half waveplate followed by two quarter waveplates, based on reference [43].
@@ II.3 Cutoff time
In the process of RSP, some qubits will have been prepared and will be sitting in the memory, waiting for the preparation of the remaining qubits. Depending on both server efficiency fibre losses, remotely preparing a qubit might take many tries, meaning that the qubits may have to reside in the memory for a long time. The longer the qubits are in memory, the more they decohere (details in section IV.1). For a lower error probability, we want to limit the time qubits spend in memory, for which we include a cutoff time.
Choosing a good cutoff time is a trade-off between rate and fidelity: a shorter cutoff time gives higher quality qubits as they have suffered from less decoherence, yet more qubits will be discarded, leading to a lower rate. However, since this work only targets an error probability (which depends on the fidelity of the remotely prepared qubits) and does not consider rate, no optimisation can be performed on this aspect. See reference [44] for examples of optimising this. Instead, we choose a fixed cutoff of half the coherence time.
When any of the qubits in memory have been at the server for longer than the cutoff time, they are discarded and the client is notified. When the server has prepared five qubits, no more qubits will be discarded (i.e., no cutoff is imposed after RSP).
@@ II.4 Graph formation
The client instructs the server to arrange the qubits in the linear graph: edges should be made between qubits $(0,1),(1,2),(2,3)$ and $(3,4)$.
We assume that the ion qubits can be individually addressed for rotations $R_{x},R_{y},R_{z}$. This ability is hardware dependent, it is possible in, for example, references [45, 46, 47], each of which uses a different mechanism to implement individual addressing. We note that it is sufficient to implement individual addressing for rotations around just one axis together with collective global rotations around an arbitrary axis in the orthogonal plane [48]. We assume the bichromatic gate $R_{XX}(\theta)=\exp(-i\theta/2\sum_{k<l}\sigma_{x}^{(k)}\sigma_{x}^{(l)})$, also known as the Mølmer-Sørensen gate [49, 50], to be the native entangling operation. We use Qiskit [51] to construct a $CZ$ gate using the operations available for the ion trap (figure 3).
In principle, the server does not need to wait for all qubits to be prepared before it starts applying the $CZ$ operations, but this is done in the simulation for simplicity. The time it takes to prepare the graph state (around 150 $\mu$s) is short compared to the expected time it takes for the qubits to be remotely prepared (on the order of tens of milliseconds [^1] The average time it takes to remotely prepare a qubit is calculated as time it takes to perform one attempt divided by the average probability of a photon arriving at the client. The average time is calculated as the distance back and forth (send qubit to client and wait for confirmation) divided by the speed of light in fibre: 2x50 km / ($c/n$) $\approx$ 0.50 ms, with $c$ being the speed of light, and $n\approx 1.5$ being the index of refection of standard telecommunication optical fibre. The arrival probability is calculated as the server efficiency (given in table 2) times the probability of photon transmitting through the fibre: 0.1325 (from table 2) $\times$ (50km $\times$ 0.2 dB/km)=0.1. This gives an average probability of photon arriving of 0.01325. Thus, it takes on average 0.50/0.01325=37.8 ms to remotely prepare a qubit.).
@@ II.5 Server measurements and output
Once the graph-state formation is complete, the client instructs the server to measure in ±$\pm_{\delta_{v}}$-bases defined by [32]
@@
Here, $\theta_{v}$ defines the basis used by the client to perform RSP, and $m_{v}$ is the outcome of the RSP measurement of the photonic qubit entangled with ion qubit $v$. The random bit $r_{v}$ is generated by the client to ensure the measurement outcomes appear random. The angles at which the dummy qubits are measured is irrelevant to the client; they just need to be random, so that the server will not be able to identify the qubit as a dummy (which would reveal that the client is executing a test round).
The server communicates the outcome of measuring qubit $v$ to the client as $b_{v}$. For all trap qubits, the client checks if $b_{v}=r_{v}$. The test round succeeds if this is true for all trap qubits. If any of the trap qubits do not satisfy $b_{v}=r_{v}$, the round fails. The round failing or succeeding is the output for each of the test rounds.